# A Unified View of Optimal Kernel Hypothesis Testing

### v1

@@ 4.3 Uniform power against alternatives separated in L2 metric
We now report power guarantees in terms of uniform separation rates with respect to the $L^{2}$-norm of the difference in densities for the two-sample problem under the standard, efficient and private testing frameworks. We consider translation-invariant kernels and impose a Sobolev smoothness requirement on the difference in densities $p-q$.
@@
The uniform separation rate of the MMD test with an optimal bandwidth depending on the unknown Sobolev smoothness $s$ is (Section A.7 and Schrab et al., 2023, Corollary 7)
$\|p-q\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N}\right)^{2s/(4s+d)}$ (21)
which is minimax optimal (Schrab et al., 2023, Appendix D). If the difference in densities is not smooth (i.e., $s\to 0$), the rate becomes constant and power cannot be guaranteed against any alternative. If $p-q$ is very smooth (i.e., $s\to\infty$), then the uniform separation rates simply becomes of order $N^{-1/2}$. We also remark that the rate deteriorates as the dimension $d$ increases.
Since the kernel bandwidth for the optimal test above depends on the unknown Sobolev smoothness, it cannot be implemented in practice. By aggregating over various kernel bandwidths (all independent of the unknown Sobolev smoothness $s$) with multiple testing, we can construct an implementable test which achieves the uniform separation rate (Section A.9 and Schrab et al., 2023, Corollary 10)
$\|p-q\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N/\log(\log(N))}\right)^{2s/(4s+d)}.$ (22)
@@
The efficient test, with an MMD estimator computable in time $\mathcal{O}\!\left(|{\mathcal{D}}|\right)$ and with optimal kernel bandwidth (depending on the unknown Soboleve smoothness $s$), controls the type II error by $\beta$ when (Section A.8 and Schrab et al., 2022b, Theorem 2)
$\|p-q\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{|{\mathcal{D}}|/N}\right)^{2s/(4s+d)}.$ (23)
When the complete statistics are used (i.e., $|{\mathcal{D}}|\asymp N^{2}$) the rate is minimax optimal as in the standard testing framework above. As the complexity $|{\mathcal{D}}|$ decreases from $N^{2}$ to $N$, the uniform separation rate gradually deteriorates, until it no longer converges to zero. This means that the result of Equation 23 does not guarantee power for the linear-time MMD tests against any alternative. However, due to the dependence of the bandwidth on the unknown Sobolev smoothness, this test cannot be implemented in practice.
Multiple testing via aggregation over a well-chosen collection of bandwidths (Schrab et al., 2022b, Theorem 2.ii), which is independent of the unknown Sobolev smoothness $s$, results in the uniform sepration rate (Section A.9 and Schrab et al., 2022b, Theorem 3)
$\|p-q\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\big{(}|{\mathcal{D}}|/N\big{)}\big{/}\log\!\big{(}\!\log(|{\mathcal{D}}|/N)\big{)}}\right)^{2s/(4s+d)}$ (24)
which is the same as the rate of Equation 23 up to an iterated logarithmic term $\log|{\mathcal{K}}|$ where $|{\mathcal{K}}|\asymp\log(|{\mathcal{D}}|/N)$ (Schrab et al., 2022b, Theorem 2.ii).
@@
The dpMMD test (Kim and Schrab, 2023, Algorithm 1 and Section 4) which is $(\varepsilon,\delta)$-differentially private achieves different uniform separation rates depending on the privacy regime (Section A.11 and Kim and Schrab, 2023, Theorem 9). Let $\xi=\varepsilon+\log\!\big{(}{1}/({1-\delta})\big{)}$. In the low privacy regime with $\xi\gtrsim N^{-(2s-d/2)/(4s+d)}$, power is guaranteed when
$\|p-q\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ N^{-2s/(4s+d)}$ (25)
@@
$\|p-q\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ (N^{{3}/{2}}\xi)^{-s/(2s+d)},$ (26)
@@
$\|p-q\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \left(N\xi\right)^{-2s/(2s+d)}.$ (27)
These rates and privacy regimes are (not yet) guaranteed to be minimax optimal, deriving matching $L^{2}$ lower bounds for differentially private testing remains an open problem, which is left for future work. These results are derived with a logarithmic dependence in $\alpha$, as mentioned in Section A.11, we believe that obtaining a logarithmic dependence in $\beta$ is also possible as in the standard and efficient testing setting.
@@ 5 Independence testing
We now consider the non-parametric independence testing problem. In Section 5.1, we formally introduce the framework, its two null simulation methods (permutations and wild bootstrap) leading to a well-calibrated non-asymptotic test. In Section 5.2 and Section 5.3, we provide power guarantees in terms of HSIC and $L^{2}$ Sobolev uniform separation rates, respectively, for independence testing under four different settings. We refer the reader to Schrab (2025b, Section 2.2) for a detailed introduction to the Hilbert–Schmidt Independence Criterion (HSIC).
@@ 5.1 Framework, bootstrap and level
Given paired samples $(X_{1},Y_{1}),\dots,(X_{N},Y_{N})$ drawn i.i.d. from a joint distribution $P_{XY}$, the aim is to test whether the first and second components of the pairs are independent, that is, ${\mathcal{H}}_{0}\colon P_{XY}=P_{X}\otimes P_{Y}$, or dependent, i.e., ${\mathcal{H}}_{1}\colon P_{XY}\neq P_{X}\otimes P_{Y}$. Mirroring the notation of Section 1 leads to defining ${\mathcal{P}}$ as the space of all joint distributions, ${\mathcal{P}}_{0}$ as the subspace of all products of marginals $\{P_{XY}\in{\mathcal{P}}:P_{XY}=P_{X}\otimes P_{Y}\}$, and ${\mathcal{P}}_{1}$ as $\{P_{XY}\in{\mathcal{P}}:P_{XY}=P_{X}\otimes P_{Y}\}$.
@@
Consider the two-sample testing problem where we are given i.i.d. samples $X_{1},\dots,X_{m}$ from a distribution $P$, and i.i.d. samples $Y_{1},\dots,Y_{n}$ from a distribution $Q$, all independent from each other, and we are interested in testing whether $P=Q$. To frame this as an independence problem, we need to construct paired samples with dependence when $P\neq Q$, and no dependence when $P=Q$. Consider the paired samples $(X_{1},1),\dots,(X_{m},1),(Y_{1},-1),\dots,(Y_{n},-1)$ where the second component of the pairs is an indicator of which sample the data point belongs to. When $P=Q$, the two components of the pairs are independent, and when $P\neq Q$, the two components are dependent. Hence, performing an independence test on these paired samples solves the two-sample problem. Computing the HSIC $V$-statistic on such paired samples with an indicator kernel for the labels is equivalent to computing a scaled MMD $V$-statistic on the original two-sample data (see Schrab, 2025b, Equation 31), and the same holds for the HSIC and MMD metrics (see paragraph ‘MMD as an HSIC’ in Schrab, 2025b, Section 2.2). If noise is required in the second component of the pairs, it can simply be added, for example, by adding i.i.d. uniform noise on $[-\nicefrac{{1}}{{2}},\nicefrac{{1}}{{2}}]$, in which case the same equivalence results between MMD and HSIC hold provided a kernel $k^{\mathcal{Y}}(y,y^{\prime})=\bm{1}(|y-y^{\prime}|\leq 1)$ is used for the labels.
@@
Independence testing can also be framed as testing for exchangeability (Section 1). To permute paired samples, only the elements in the second component of the pairs are permuted. Independent paired samples (i.e., null hypothesis) are exactly the ones which are exchangeable. Indeed, clearly, the independent paired samples are exchangeable since there is no dependence between the two components of the pairs. Under the alternative, permuting the elements of the second component breaks the dependence, so the paired samples are not exchangeable as the original and permuted paired samples are not identically distributed. Therefore, independence testing corresponds to testing for exchangeability, and the independence null hypothesis can be simulated using permutations.
@@
As presented above, given a permutation $\pi$ of $\{1,\dots,N\}$, permuting the original paired samples $\mathbb{Z}_{N}=((X_{i},Y_{i}))_{i=1}^{N}$ with respect to $\pi$ results in $((X_{i},Y_{\pi(i)}))_{i=1}^{N}$. For any statistic function $T$, the test statistic is $T(\mathbb{Z}_{N})$, and a permuted statistic can be computed as $T(\mathbb{Z}_{N}^{\pi})$ for any permutation $\pi$. Following Section 1.2, a test can be constructed using randomly sampled permutations, it achieves type I error control at the prescribed level $\alpha$ non-asymptotically (Romano and Wolf, 2005a, Lemma 1, see also Kim and Schrab, 2023, Lemma 15).
@@
When using a one-sample second-order statistic (see Schrab, 2025b, Section 3), the null can be simulated asymptotically with wild bootstrapped statistics ${|{\mathcal{D}}|}^{-1}\sum_{(i,j)\in{\mathcal{D}}}\varepsilon_{i}\varepsilon_{j}h(X_{i},X_{j})$ using i.i.d. Rademacher variables $\varepsilon_{1},\dots,\varepsilon_{n}$. Relying on these, a test controlling the level asymptotically (Chwialkowski et al., 2014, Theorem 1) can be constructed following the procedure described in Section 1.2. When using the HSIC estimator of Schrab (2025b, Equation 25) with $h^{\operatorname{HSIC}}_{{k^{\mathcal{X}}}\!,{k^{\mathcal{Y}}}}$ as in Schrab (2025b, Equation 23), computing a wild bootstrapped statistic corresponds to computing a permuted statistic for some specific permutation allowed to swap $i$ with $i+N/2$ for $i=1,\dots,N/2$, for $N$ even (see Schrab et al., 2022b, Appendix F.1 for details). Leveraging this fact, non-asymptotic level control can be guaranteed for this HSIC wild bootstrap test. Discussions regarding efficient implementations of such tests, as well as guidance on when to use each bootstrapping methods, are provided in Section 1.2. See also Pogodin et al. (2024, Theorem 4) for wild bootstrap guarantees for quantities closely related to HSIC.
@@
The permutation-based HSIC, dpHSIC and dcHSIC tests (Section 3), as well as the wild bootstrap HSIC test, all control the probability of type I error at $\alpha$ at every sample size as desired (Albert et al., 2022, Proposition 1; Kim and Schrab, 2023, Theorem 6; Schrab and Kim, 2025, Lemmas 1 and 4). This non-asympotic level is preserved when using efficient estimators (Schrab et al., 2022b, Proposition 1), as well as when using adaptivity over kernels, either via pooling (properties of the permutation method, Romano and Wolf, 2005a, Lemma 1, combined with Albert et al. 2022, Proposition 1; see also discussion around Biggs et al., 2023, Theorem 1) or via aggregation (Albert et al., 2022, Section 3.1).
@@
The power of the HSIC tests is greatly affected by the choice of the two kernels. This issue of kernel selection can be addressed either via aggregation (Section 2) or via kernel pooling (Schrab, 2025b, Section 4). In practice, we recommend using the HSICAgg test (Albert et al., 2022) and the normalised HSICFuse test.
@@ 5.2 Uniform power against alternatives separated in HSIC metric
The HSIC-based independence test with fixed kernels $k$ and $\ell$ achieves high power if (Section A.1 and Kim and Schrab, 2023, Theorem 12)
$\mathrm{HSIC}_{k,\ell}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}}$ (28)
which is minimax optimal (Kim and Schrab, 2023, Theorem 13). Using mean kernel pooling (Schrab, 2025b, Section 4) results in the uniform separation rate (Section A.4)
$\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\leavevmode\nobreak\ \underset{\ell\in{\mathcal{L}}}{\mathrm{mean}}\leavevmode\nobreak\ \mathrm{HSIC}_{k,\ell}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}}.$ (29)
Relying on unnormalised fuse/max kernel pooling for the HSIC test results in an additional logarithmic term in the size of the product kernel collection, that is (Section A.4)
$\max_{k\in{\mathcal{K}}}\,\max_{\ell\in{\mathcal{L}}}\,\mathrm{HSIC}_{k,\ell}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}||{\mathcal{L}}|)\big{\}}}{N}}$ (30)
where the fusing parameter is assumed to satisfy $\nu\geq\max(N,\log(|{\mathcal{K}}||{\mathcal{L}}|))$, and where a typical choice is $|{\mathcal{K}}|=|{\mathcal{L}}|=\log(N)$ (e.g., Schrab et al., 2023, Corollary 10). [^7] The fuse extension from the MMD two-sample framework to the HSIC independence one has brilliantly been conducted by Ren (Michael) Guangyo as part of his UCL MSc Machine Learning Project supervised by Antonin Schrab and Arthur Gretton. As in the MMD case, for other testing constraints, we present only the results for max and fuse kernel pooling (not necessarily always mentioning the assumption on $\nu$), but the results for mean pooling hold with the same rate as for fixed kernel, similarly to Equation 29.
@@
The efficient HSIC test, with block HSIC B-statistic (Schrab, 2025b, Equation 70) consisting of $B$ blocks, and fixed kernels $k$ and $\ell$, controls the type II error by $\beta$ when (Section A.5)
$\mathrm{HSIC}_{k,\ell}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{B\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}}$ (31)
which achieves the standard minimax optimal rate of Equation 28 when the complete U-statistic is used (i.e., $B=1$). As the number of blocks $B$ is increased from $1$ to $N$, the uniform separation rate gradually slows down from $N^{-1/2}$ to $N^{0}$ (i.e., no longer converging to zero). The unnormalised pooled fuse/max block HSIC test (Schrab, 2025b, Section 4) has uniform separation rate (Section A.4)
$\max_{k\in{\mathcal{K}}}\,\max_{\ell\in{\mathcal{L}}}\,\mathrm{HSIC}_{k,\ell}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{B\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}||{\mathcal{L}}|)\big{\}}}{N}}$ (32)
@@
The $(\varepsilon,\delta)$-differentially private dpHSIC test (Kim and Schrab, 2023, Algorithm 1 and Appendix B.5) with fixed kernels $k$ and $\ell$ is powerful when (Section A.2 and Kim and Schrab, 2023, Theorem 12)
$\mathrm{HSIC}_{k,\ell}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \max\Biggl{\{}\sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}},\,\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N\xi}\Biggr{\}}.$ (33)
@@
For differential privacy, kernel pooling is rendered difficult due to the fact that the privatisation noise would scale with the number of kernels, unless proved otherwise (see discussion below Equation 18 for details).
@@
Being robust against corruption of up to $r$ samples, the dcHSIC (Schrab and Kim, 2025, Algorithm 1 and Section 4) and dpHSIC (Schrab and Kim, 2025, Algorithm 2, Section 5, Appendix E) tests with fixed kernels $k$ and $\ell$ have uniform separation rate (Section A.3 and Schrab and Kim, 2025, Theorem 2.i and 4)
$\mathrm{HSIC}_{k,\ell}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \max\biggl{\{}\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta)\}}{N}},\ \frac{r}{N}\biggr{\}}$ (34)

### v2

@@ 4.3 Uniform power against alternatives separated in L2 metric
We now report power guarantees in terms of uniform separation rates with respect to the $L^{2}$-norm of the difference in densities for the two-sample problem under the standard, efficient and private testing frameworks. We consider translation-invariant kernels and impose a Sobolev smoothness requirement on the difference in densities pqp-q.
@@
The uniform separation rate of the MMD test with an optimal bandwidth depending on the unknown Sobolev smoothness $s$ is (Section˜A.7 and Schrab et al., 2023, Corollary 7)
$\|p-q\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N}\right)^{2s/(4s+d)}$ (21)
which is minimax optimal (Schrab et al., 2023, Appendix D). If the difference in densities is not smooth (i.e., $s\to 0$), the rate becomes constant and power cannot be guaranteed against any alternative. If pqp-q is very smooth (i.e., $s\to\infty$), then the uniform separation rates simply becomes of order $N^{-1/2}$. We also remark that the rate deteriorates as the dimension $d$ increases.
Since the kernel bandwidth for the optimal test above depends on the unknown Sobolev smoothness, it cannot be implemented in practice. By aggregating over various kernel bandwidths (all independent of the unknown Sobolev smoothness $s$) with multiple testing, we can construct an implementable test which achieves the uniform separation rate (Section˜A.9 and Schrab et al., 2023, Corollary 10)
$\|p-q\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N/\log(\log(N))}\right)^{2s/(4s+d)}.$ (22)
@@
The efficient test, with an MMD estimator computable in time $\mathcal{O}\!\left(|{\mathcal{D}}|\right)$ and with optimal kernel bandwidth (depending on the unknown Soboleve smoothness $s$), controls the type II error by $\beta$ when (Section˜A.8 and Schrab et al., 2022b, Theorem 2)
$\|p-q\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{|{\mathcal{D}}|/N}\right)^{2s/(4s+d)}.$ (23)
When the complete statistics are used (i.e., $|{\mathcal{D}}|\asymp N^{2}$) the rate is minimax optimal as in the standard testing framework above. As the complexity $|{\mathcal{D}}|$ decreases from $N^{2}$ to $N$, the uniform separation rate gradually deteriorates, until it no longer converges to zero. This means that the result of Equation˜23 does not guarantee power for the linear-time MMD tests against any alternative. However, due to the dependence of the bandwidth on the unknown Sobolev smoothness, this test cannot be implemented in practice.
Multiple testing via aggregation over a well-chosen collection of bandwidths (Schrab et al., 2022b, Theorem 2.ii), which is independent of the unknown Sobolev smoothness $s$, results in the uniform sepration rate (Section˜A.9 and Schrab et al., 2022b, Theorem 3)
$\|p-q\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\big(|{\mathcal{D}}|/N\big)\big/\log\!\big(\!\log(|{\mathcal{D}}|/N)\big)}\right)^{2s/(4s+d)}$ (24)
which is the same as the rate of Equation˜23 up to an iterated logarithmic term $\log|{\mathcal{K}}|$ where $|{\mathcal{K}}|\asymp\log(|{\mathcal{D}}|/N)$ (Schrab et al., 2022b, Theorem 2.ii).
@@
The dpMMD test (Kim and Schrab, 2023, Algorithm 1 and Section 4) which is $(\varepsilon,\delta)$-differentially private achieves different uniform separation rates depending on the privacy regime (Section˜A.11 and Kim and Schrab, 2023, Theorem 9). Let $\xi=\varepsilon+\log\!\big({1}/({1-\delta})\big)$. In the low privacy regime with $\xi\gtrsim N^{-(2s-d/2)/(4s+d)}$, power is guaranteed when
$\|p-q\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ N^{-2s/(4s+d)}$ (25)
@@
$\|p-q\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ (N^{{3}/{2}}\xi)^{-s/(2s+d)},$ (26)
@@
$\|p-q\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \left(N\xi\right)^{-2s/(2s+d)}.$ (27)
These rates and privacy regimes are (not yet) guaranteed to be minimax optimal, deriving matching $L^{2}$ lower bounds for differentially private testing remains an open problem, which is left for future work. These results are derived with a logarithmic dependence in $\alpha$, as mentioned in Section˜A.11, we believe that obtaining a logarithmic dependence in $\beta$ is also possible as in the standard and efficient testing setting.
@@ 5 Independence testing
We now consider the non-parametric independence testing problem. In Section˜5.1, we formally introduce the framework, its two null simulation methods (permutations and wild bootstrap) leading to a well-calibrated non-asymptotic test. In Section˜5.2 and Section˜5.3, we provide power guarantees in terms of HSIC and $L^{2}$ Sobolev uniform separation rates, respectively, for independence testing under four different settings. We refer the reader to Schrab (2025b, Section 2.2) for a detailed introduction to the Hilbert–Schmidt Independence Criterion (HSIC).
@@ 5.1 Framework, bootstrap and level
Given paired samples $(X_{1},Y_{1}),\dots,(X_{N},Y_{N})$ drawn i.i.d. from a joint distribution $P_{XY}$, the aim is to test whether the first and second components of the pairs are independent, that is, ${\mathcal{H}}_{0}\colon P_{XY}=P_{X}\otimes P_{Y}$, or dependent, i.e., ${\mathcal{H}}_{1}\colon P_{XY}\neq P_{X}\otimes P_{Y}$. Mirroring the notation of Section˜1 leads to defining $\mathcal{P}$} as the space of all joint distributions, ${\mathcal{P}}_{0}$ as the subspace of all products of marginals $\{P_{XY}\in{\mathcal{P}}:P_{XY}=P_{X}\otimes P_{Y}\}$, and ${\mathcal{P}}_{1}$ as $\{P_{XY}\in{\mathcal{P}}:P_{XY}=P_{X}\otimes P_{Y}\}$.
@@
Consider the two-sample testing problem where we are given i.i.d. samples $X_{1},\dots,X_{m}$ from a distribution $P$, and i.i.d. samples $Y_{1},\dots,Y_{n}$ from a distribution $Q$, all independent from each other, and we are interested in testing whether $P=Q$. To frame this as an independence problem, we need to construct paired samples with dependence when $P\neq Q$, and no dependence when $P=Q$. Consider the paired samples $(X_{1},1),\dots,(X_{m},1),(Y_{1},-1),\dots,(Y_{n},-1)$ where the second component of the pairs is an indicator of which sample the data point belongs to. When $P=Q$, the two components of the pairs are independent, and when $P\neq Q$, the two components are dependent. Hence, performing an independence test on these paired samples solves the two-sample problem. Computing the HSIC $V$-statistic on such paired samples with an indicator kernel for the labels is equivalent to computing a scaled MMD $V$-statistic on the original two-sample data (see Schrab, 2025b, Equation 31), and the same holds for the HSIC and MMD metrics (see paragraph ‘MMD as an HSIC’ in Schrab, 2025b, Section 2.2). If noise is required in the second component of the pairs, it can simply be added, for example, by adding i.i.d. uniform noise on $\nicefrac{{1}}{{2}},\nicefrac{{1}}{{2}}]$, in which case the same equivalence results between MMD and HSIC hold provided a kernel $k^{\mathcal{Y}}(y,y^{\prime})=\bm{1}(|y-y^{\prime}|\leq 1)$ is used for the labels.
@@
Independence testing can also be framed as testing for exchangeability (Section˜1). To permute paired samples, only the elements in the second component of the pairs are permuted. Independent paired samples (i.e., null hypothesis) are exactly the ones which are exchangeable. Indeed, clearly, the independent paired samples are exchangeable since there is no dependence between the two components of the pairs. Under the alternative, permuting the elements of the second component breaks the dependence, so the paired samples are not exchangeable as the original and permuted paired samples are not identically distributed. Therefore, independence testing corresponds to testing for exchangeability, and the independence null hypothesis can be simulated using permutations.
@@
As presented above, given a permutation $\pi$ of $\{1,\dots,N\}$, permuting the original paired samples $\mathbb{Z}_{N}=((X_{i},Y_{i}))_{i=1}^{N}$ with respect to $\pi$ results in $((X_{i},Y_{\pi(i)}))_{i=1}^{N}$. For any statistic function $T$, the test statistic is $T(\mathbb{Z}_{N})$, and a permuted statistic can be computed as $T(\mathbb{Z}_{N}^{\pi})$ for any permutation $\pi$. Following Section˜1.2, a test can be constructed using randomly sampled permutations, it achieves type I error control at the prescribed level $\alpha$ non-asymptotically (Romano and Wolf, 2005a, Lemma 1, see also Kim and Schrab, 2023, Lemma 15).
@@
When using a one-sample second-order statistic (see Schrab, 2025b, Section 3), the null can be simulated asymptotically with wild bootstrapped statistics ${|{\mathcal{D}}|}^{-1}\sum_{(i,j)\in{\mathcal{D}}}\varepsilon_{i}\varepsilon_{j}h(X_{i},X_{j})$ using i.i.d. Rademacher variables $\varepsilon_{1},\dots,\varepsilon_{n}$. Relying on these, a test controlling the level asymptotically (Chwialkowski et al., 2014, Theorem 1) can be constructed following the procedure described in Section˜1.2. When using the HSIC estimator of Schrab (2025b, Equation 25) with $h^{\operatorname{HSIC}}_{{k^{\mathcal{X}}}\!,{k^{\mathcal{Y}}}}$ as in Schrab (2025b, Equation 23), computing a wild bootstrapped statistic corresponds to computing a permuted statistic for some specific permutation allowed to swap $i$ with $i+N/2$ for $i=1,\dots,N/2$, for $N$ even (see Schrab et al., 2022b, Appendix F.1 for details). Leveraging this fact, non-asymptotic level control can be guaranteed for this HSIC wild bootstrap test. Discussions regarding efficient implementations of such tests, as well as guidance on when to use each bootstrapping methods, are provided in Section˜1.2. See also Pogodin et al. (2024, Theorem 4) for wild bootstrap guarantees for quantities closely related to HSIC.
@@
The permutation-based HSIC, dpHSIC and dcHSIC tests (Section˜3), as well as the wild bootstrap HSIC test, all control the probability of type I error at $\alpha$ at every sample size as desired (Albert et al., 2022, Proposition 1; Kim and Schrab, 2023, Theorem 6; Schrab and Kim, 2025, Lemmas 1 and 4). This non-asympotic level is preserved when using efficient estimators (Schrab et al., 2022b, Proposition 1), as well as when using adaptivity over kernels, either via pooling (properties of the permutation method, Romano and Wolf, 2005a, Lemma 1, combined with Albert et al. 2022, Proposition 1; see also discussion around Biggs et al., 2023, Theorem 1) or via aggregation (Albert et al., 2022, Section 3.1).
@@
The power of the HSIC tests is greatly affected by the choice of the two kernels. This issue of kernel selection can be addressed either via aggregation (Section˜2) or via kernel pooling (Schrab, 2025b, Section 4). In practice, we recommend using the HSICAgg test (Albert et al., 2022) and the normalised HSICFuse test.
@@ 5.2 Uniform power against alternatives separated in HSIC metric
The HSIC-based independence test with fixed kernels $k$ and $\ell$ achieves high power if (Section˜A.1 and Kim and Schrab, 2023, Theorem 12)
$\mathrm{HSIC}_{k,\ell}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}}$ (28)
which is minimax optimal (Kim and Schrab, 2023, Theorem 13). Using mean kernel pooling (Schrab, 2025b, Section 4) results in the uniform separation rate (Section˜A.4)
$\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\penalty 10000\ \underset{\ell\in{\mathcal{L}}}{\mathrm{mean}}\penalty 10000\ \mathrm{HSIC}_{k,\ell}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}}.$ (29)
Relying on unnormalised fuse/max kernel pooling for the HSIC test results in an additional logarithmic term in the size of the product kernel collection, that is (Section˜A.4)
$\max_{k\in{\mathcal{K}}}\,\max_{\ell\in{\mathcal{L}}}\,\mathrm{HSIC}_{k,\ell}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}||{\mathcal{L}}|)\big\}}{N}}$ (30)
where the fusing parameter is assumed to satisfy $\nu\geq\max(N,\log(|{\mathcal{K}}||{\mathcal{L}}|))$, and where a typical choice is $|{\mathcal{K}}|=|{\mathcal{L}}|=\log(N)$ (e.g., Schrab et al., 2023, Corollary 10). [^7] The fuse extension from the MMD two-sample framework to the HSIC independence one has brilliantly been conducted by Ren (Michael) Guangyo as part of his UCL MSc Machine Learning Project supervised by Antonin Schrab and Arthur Gretton. As in the MMD case, for other testing constraints, we present only the results for max and fuse kernel pooling (not necessarily always mentioning the assumption on $\nu$), but the results for mean pooling hold with the same rate as for fixed kernel, similarly to Equation˜29.
@@
The efficient HSIC test, with block HSIC B-statistic (Schrab, 2025b, Equation 70) consisting of $B$ blocks, and fixed kernels $k$ and $\ell$, controls the type II error by $\beta$ when (Section˜A.5)
$\mathrm{HSIC}_{k,\ell}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{B\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}}$ (31)
which achieves the standard minimax optimal rate of Equation˜28 when the complete U-statistic is used (i.e., $B=1$). As the number of blocks $B$ is increased from $1$ to $N$, the uniform separation rate gradually slows down from $N^{-1/2}$ to $N^{0}$ (i.e., no longer converging to zero). The unnormalised pooled fuse/max block HSIC test (Schrab, 2025b, Section 4) has uniform separation rate (Section˜A.4)
$\max_{k\in{\mathcal{K}}}\,\max_{\ell\in{\mathcal{L}}}\,\mathrm{HSIC}_{k,\ell}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{B\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}||{\mathcal{L}}|)\big\}}{N}}$ (32)
@@
The $(\varepsilon,\delta)$-differentially private dpHSIC test (Kim and Schrab, 2023, Algorithm 1 and Appendix B.5) with fixed kernels $k$ and $\ell$ is powerful when (Section˜A.2 and Kim and Schrab, 2023, Theorem 12)
$\mathrm{HSIC}_{k,\ell}\penalty 10000\ \gtrsim\penalty 10000\ \max\Biggl\{\sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}},\,\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N\xi}\Biggr\}.$ (33)
@@
For differential privacy, kernel pooling is rendered difficult due to the fact that the privatisation noise would scale with the number of kernels, unless proved otherwise (see discussion below Equation˜18 for details).
@@
Being robust against corruption of up to $r$ samples, the dcHSIC (Schrab and Kim, 2025, Algorithm 1 and Section 4) and dpHSIC (Schrab and Kim, 2025, Algorithm 2, Section 5, Appendix E) tests with fixed kernels $k$ and $\ell$ have uniform separation rate (Section˜A.3 and Schrab and Kim, 2025, Theorem 2.i and 4)
$\mathrm{HSIC}_{k,\ell}\penalty 10000\ \gtrsim\penalty 10000\ \max\biggl\{\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta)\}}{N}},\ \frac{r}{N}\biggr\}$ (34)